# 3-manifolds with universal links are spherical

### v1

@@ Abstract.
We determine which orientable $3$-manifolds are dominated by the $3$-sphere. This characterization implies that the closed $3$-manifolds that admit a universal link are spherical. Moreover, we show examples of complement universal links that are not universal.
@@ 1. Introduction
Hilden and Montesinos proved that every orientable $3$-manifold is a $3$-fold branched covering of the $3$-sphere $\mathbb{S}^{3}$. With this result as a starting point, Thurston proved the striking fact that there exists a link $L\subset\mathbb{S}^{3}$ such that every orientable $3$-manifold is a branched covering of $\mathbb{S}^{3}$ with branching set exactly $L$. Thurston called $L$ a universal link. In this paper, we characterize the existence of universal links for closed $3$-manifolds:
@@
A closed $3$-manifold admits a universal link if and only if it is spherical.
This result is a consequence of a more general result in non-zero degree maps, where branched covers between orientable closed manifolds are a particular case. We define a reflexive and transitive relation between orientable closed manifolds by saying that $N$ dominates $M$ if there is a non-zero degree map from $N$ to $M$ (see Section 4). The existence of universal links implies that all orientable $3$-manifolds dominate the $3$-sphere. By definition, all spherical $3$-manifolds are dominated by the $3$-sphere, but actually these are the only $3$-manifolds that satisfy this property:
Theorem 4.1.
The top elements in the set of closed, orientable $3$-manifolds, under the domination relation, are exactly the spherical $3$-manifolds.
We give two proofs of Theorem 4.1, one using the residual finiteness for compact $3$-manifolds [Hempel:ResidualFnite], and a second using [KotschickNeofytidis:DominationByCircleBunles, Theorem 3] due to Kotschick and Neofytidis, which allows us to construct a finite cover from a connected sum of $\mathbb{S}^{2}\times\mathbb{S}^{1}$ to manifolds without aspherical prime factors in their prime decomposition.
It is worth mentioning that Edmonds proved in [Edmonds:Pi_1SurjectiveMapsBranchedCovers] that every map between closed, orientable 3-manifolds with a degree greater than $2$ is homotopic to a branched cover, which explains a reciprocal relation between non-zero degree maps and branched covers.
Finally, we notice that Thurston provided a more general definition of a universal link in [Thurston:UniversalLinks], by calling a link $L$ in an arbitrary orientable closed $3$-manifold $M$ complement universal if there exists a link in any orientable closed $3$-manifold whose complement is virtually homeomorphic to $M{\smallsetminus}L$ (see Section 5). We indeed showed that these two definitions are not equivalent:
@@
There exists a universal link in infinitely many $3$-manifolds which is not a complement universal link.
@@ 1.1. Acknowledgements
We thank Mario Eudave for pointing us toward helpful references.
@@ 2. Universal links in spherical $3$-manifolds
Theorem 2.4.
Let $M$ be a spherical 3-manifold; then $M$ has universal knots and links.
Proof.
We address the proof of the existence of universal links; however, the same can be applied to knots.
[…]
In the following section, we explore the obstruction of the existence of branched covering maps between the $3$-sphere and an orientable $3$-manifold in a wider family of maps, which are the ones with non-zero degree.
3. Transfer maps
The subject of transfer maps on covering maps is well-known. But its existence can be extended to non-zero degree maps. The results presented in this section are probably folklore knowledge among experts. However, it might not be easy to extract from the literature. Although we could track the first moment this argument was used in the proof of [BeresteinEdmonds:BranchedCovers, Theorem 7.3]. For this reason, and for the sake of being more self-contained, we provide the details here.
Definition 3.1.
Given a map $f:M^{n}\to N^{n}$ between orientable $n$-manifolds, we define the transfer of
$f_{\#}:H_{k}(M)\to H_{k}(N)$
as the map
$\tau_{f}=PD_{M}\circ f^{\#}\circ PD_{N}^{-1}:H_{k}(N)\to H_{k}(M),$
where $PD_{*}$ denotes the Poincaré duality isomorphism, and $f_{\#},f^{\#}$ are the induced maps of $f$ in homology and cohomology, respectively.
Notice that $\tau_{f}$ is the map from $H_{k}(N)$ to $H_{k}(M)$ obtained by following the commutative diagram below, without passing through $f_{\#}$ directly:
$H_{k}(M)$$H_{k}(N)$$H^{n-k}(M)$$H^{n-k}(N)$$\scriptstyle{f_{\#}}$$\scriptstyle{PD_{M}}$$\scriptstyle{f^{\#}}$$\scriptstyle{PD_{N}}$
We now prove the following result, where $\deg(f)$ denotes the degree of $f$. Recall that this degree is defined for any map between orientable manifolds $f:M^{n}\to N^{n}$ as the integer $d$ satisfying
$f_{\#}([M])=d\cdot[N]\in H_{n}(N),$
where $M]\in H_{n}(M)$ and $N]\in H_{n}(N)$ are the fundamental classes of $M$ and $N$, respectively.
Lemma 3.2.
$f_{\#}\circ\tau_{f}([\alpha])=\deg(f)\cdot[\alpha]$.
Proof.
This follows directly from the definition $\tau_{f}:=PD_{M}\circ f^{\#}\circ PD_{N}^{-1}$.
Since $PD_{N}$ is an isomorphism, there exists a unique class $a]\in H^{n-k}(N)$ such that
$PD_{N}([a])=[\alpha]\quad\text{i.e.,}\quad[a]\smallfrown[N]=[\alpha].$
Then we compute:
$f_{\#}\circ\tau_{f}([\alpha])$ $=f_{\#}\circ\tau_{f}([a]\smallfrown[N])$
$=f_{\#}\circ(PD_{M}\circ f^{\#}\circ PD^{-1}_{N})([a]\smallfrown[N])$
$=f_{\#}\circ(f^{\#}([a])\smallfrown[M])$
By the naturality of the cap product, $f_{\#}(f^{\#}([a])\smallfrown[M])=[a]\smallfrown f_{\#}([M])$, and by definition, $f_{\#}([M])=\deg(f)\cdot[N]$. Therefore,
$f_{\#}\circ\tau_{f}([\alpha])$ $=([a]\smallfrown f_{\#}([M]))$
$=([a]\smallfrown deg(f)\cdot[N])$
$=\deg(f)\cdot([a]\smallfrown[N])$
$=\deg(f)\cdot[\alpha]$
Observe that if $f:M\to N$ is an $d$-sheeted branched covering map, then $\deg(f)=d$. In particular, when $M=\mathbb{S}^{3}$, this implies that $f_{\#}\circ\tau_{f}:H_{1}(N)\to H_{1}(N)$ is simultaneously the zero map and the multiplication by $d$. This yields the following result.
Lemma 3.3.
Let $N$ be a orientable $3$-manifold. If there exists a map $f:\mathbb{S}^{3}\to N$ of non-zero degree, then $H_{1}(N)$ is a finite group; equivalently, $N$ is a rational homology sphere.
Lemma 3.4.
Let $\phi:\widetilde{M}\to M$ be a finite covering map, and let $f:\mathbb{S}^{3}\to M$ be a map of non-zero degree. Then there exists a lift $\widetilde{f}:\mathbb{S}^{3}\to\widetilde{M}$ of $f$ with non-zero degree.
[…]
Proof.
In homology, we have the relation $\phi_{\#}\circ\widetilde{f}_{\#}=f_{\#}$. Taking degrees gives
$\deg(\phi)\cdot\deg(\widetilde{f})=\deg(f)\neq 0,$
so $\deg(\widetilde{f})\neq 0$. ∎
In [Smith:TransferRamifiedCoverings], L. Smith defined an $d$-fold ramified covering map, where branched coverings are a particular case, and proved a generalization of Lemma 3.2 that includes non-orientable manifolds, as follows:
@@ 3. Transfer maps
4. Top elements of closed $3$-manifolds under domination relation
The existence of a non-zero degree map, $N\rightarrow M$, defines a reflexive and transitive relation on the set of homotopy types of closed orientable manifolds. Whenever such a map exists, we say that $N$ dominates $M$ and write $N\geq M$. In this case $N$ is at least as complicated as $M$. For example, in Lemma 3.2 we saw that the induced maps in rational homology are surjective, thus, in particular, the Betti numbers of $M$ are bounded from above by those of $N$. Also, $N\geq M$ implies that the fundamental group of $N$ surjects onto a finite index subgroup of the fundamental group of $M$.
@@ 4. Top elements of closed $3$-manifolds under domination relation
The top elements in the set of closed, orientable $3$-manifolds, under the domination relation, are exactly the spherical $3$-manifolds.
As every $3$-manifold is dominated by spherical $3$-manifolds by Theorem 2.4, it is enough to show that if $M$ is dominated by $\mathbb{S}^{3}$ then $M$ has a finite fundamental group. We provide two proofs of this fact:
First proof.
As non-spherical irreducible, closed, orientable $3$-manifolds are aspherical, then there is no domination of $\mathbb{S}^{3}$ to them. So we focus on the reducible case.
On one hand, if $M$ is prime, then $M=\mathbb{S}^{1}\times\mathbb{S}^{2}$ and it can not be dominated by $\mathbb{S}^{3},$ because of Lemma 3.3.
On the other hand, if $M$ has a separating $2$-sphere $S,$ such that $M=M_{1}\sharp_{S}M_{2}$, the residual finiteness for compact $3$-manifolds [Hempel:ResidualFnite], implies that there is $K_{i}$ a proper normal subgroup of $\pi_{1}(M_{i})$ of finite index, for each $i=1,2$.
Let $\phi:N\rightarrow M$ be the finite cover induced by the kernel of the following epimorphism:
Consider $p\in S,$ $q\in\phi^{-1}(p)$ and $S_{1}$ the connected component of $\phi^{-1}(S)$ that contains $q.$
We claim that that $S_{1}$ is a non-separating $2$-sphere. Indeed, take a simple loop $x_{i}:([0,1],0)\rightarrow(M_{i},p)$ whose homotopy class belongs to $\pi_{1}(M_{i}){\smallsetminus}K_{i}.$ Notice that the homotopy class $x_{1}][x_{2}][x_{1}]^{-1}[x_{2}]^{-1}\in Ker(\phi).$ Then the corresponding loop, lifts to a loop in $N,$ denoted by $s.$ Notice that neither $x_{1}],[x_{1}][x_{2}]$ nor $x_{1}][x_{2}][x_{1}]^{-1}$ are in $Ker(\phi),$ so $s$ intersects once $S_{1}$ at $q.$ And the intersection is transversal, because $s$ starts inside $\phi^{-1}(M_{1})$ and ends inside $\phi^{-1}(M_{2}),$ which are in different sides of $S_{1}.$
After taking a regular neighborhood of $s\cup S_{1},$ we have that NN has $\mathbb{S}^{1}\times\mathbb{S}^{2}$ as a prime summands in its prime decomposition, but by Lemma 3.4, and Lemma 3.3 we have that $N$ is a rational homology sphere, which is impossible. ∎
Second proof.
By the Kneser-Milnor prime decomposition, $M$ has an essentially unique prime decomposition $M=M_{1}\sharp...\sharp M_{k}$ under the connected sum operation. Each prime summand $M_{i}$ is either aspherical, $\mathbb{S}^{1}\times\mathbb{S}^{2},$ or has a finite fundamental group.
Suppose we have a non-zero degree map $f:\mathbb{S}^{3}\rightarrow M.$ If $M$ has an aspherical summand $M_{i}$ in its prime decomposition, then by composing $f$ with the degree one map $M\rightarrow M_{i}$ that collapses the connected summands other than $M_{i},$ we obtain a dominant map $\mathbb{S}^{3}\rightarrow M_{i},$ but that is impossible because $\pi_{3}(M_{i})=\{0\}$.
So, if we have a non-zero degree map $f:\mathbb{S}^{3}\rightarrow M,$ then $M$ does not have an aspherical summand $M_{i}$ in its prime decomposition. By [KotschickNeofytidis:DominationByCircleBunles, Theorem 3] we conclude that $M$ is finitely covered by a connected sum $\sharp_{k}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ where the empty connected sum $(k=0)$ denotes the $3$-sphere $\mathbb{S}^{3}.$
Notice that if $k>0$ then we could lift $f:\mathbb{S}^{3}\rightarrow M$ to a non-zero degree map $\widetilde{f}:\mathbb{S}^{3}\rightarrow\sharp_{k}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ by Lemma 3.4, and Lemma 3.3 would imply that $\sharp_{k}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ is a rational homology sphere, which is impossible. Then $k=0$, meaning that $M$ is a spherical 33-manifold. ∎
Corollary 4.2.
A closed $3$-manifold admits universal links if and only if it is spherical.
@@
By Theorem 4.1 and Theorem 2.4 we have the case when $M$ is an orientable closed 33-manifold. For the non-orientable case, notice that $M$ non-zero first Betti number [BeresteinEdmonds:BranchedCovers, Lemma 8.6]. Then by Proposition 3.5, we have that there is not branching cover from the $3$-sphere to $M.$
@@ 5. Complement universal links that are not universal
Notice that if $L$ is a universal link, it is a complement universal link, because the restriction of a branched covering map on the complement of the branching set is a finite covering map (see Def. 2.2).
@@
There exists a complement universal link in infinitely many $3$-manifolds that is not a universal.
@@
Consider $M$, a non-spherical $3$-manifold that is homeomorphic to a Dehn surgery over a hyperbolic universal link $L$ in a spherical $3$-manifold $S.$ By Thurston’s hyperbolic Dehn filling theorem, this provides infinitely many $3$-manifolds $M.$ Moreover, for every $N$ orientable closed $3$-manifold there is a branched cover $p:(N,p^{-1}(L))\rightarrow(S,L),$ which induces a finite cover from $N{\smallsetminus}p^{-1}(L)$ to $S{\smallsetminus}L,$ and the latter is homeomorphic to the complement of the core $L^{\prime}$ of the Dehn filled tori in $M.$
@@
In conclusion, $L^{\prime}\subset M$ is a complement universal link, that is not a universal link because $M$ is non-spherical by Corollary 4.2. And by varying over the integral Dehn surgery on $L$ in $S$, we obtain infinitely many non-homeomorphic $3$-manifolds $M.$

### v3

# On the existence of universal links in three-manifolds

@@ Abstract.
We study the existence of branched coverings between closed $3$-manifolds, with emphasis on universal knots and links. We prove that the only closed $3$-manifolds that admit a universal link are spherical. Furthermore, we distinguish between universal links and complement universal links and show that these notions do not coincide in general, by exhibiting infinitely many examples of complement universal links that are not universal. Also, we prove that there is no closed aspherical $3$-manifold, such that every closed, aspherical $3$-manifold is a branched covering over it. Finally, we characterize the closed $3$-manifolds admitting branching coverings from $P^{3}\#P^{3}$, and deduce that there is no closed reducible $3$-manifold, such that every closed reducible $3$-manifold is a branched covering over it.
@@ 1. Introduction
Hilden proved in [Hilde:BranchedCoverToS3] that every orientable $3$-manifold is a $3$-fold branched covering of the $3$-sphere $\mathbb{S}^{3}$. With this result as a starting point, Thurston proved in [Thurston:UniversalLinks] the striking fact that there exists a link $L\subset\mathbb{S}^{3}$ such that every orientable $3$-manifold is a branched covering of $\mathbb{S}^{3}$ with branching set exactly $L$. Thurston called $L$ a universal link. The existence of universal links reflects geometric and group-theoretic constrains on the ambient manifold. In particular, it is natural to ask for a classification of closed $3$-manifolds that admit such links. In this paper, we answer this question:
Theorem 2.8.
A closed $3$-manifold admits universal knots and links, if and only if, it is spherical.
The necessity condition for the orientable case in Theorem 2.8, comes from the fact that given any branched cover, between orientable closed $3$-manifolds $f:N\rightarrow M$, then $f_{*}:\pi_{1}(N)\rightarrow\pi_{1}(M)$ has an image of finite index in $\pi_{1}(M),$ which was first notice by Hopf in [Hopf:ImageOfFiniteIndex, Satz VIII] (see Proposition 2.2). This means that if we have a branched cover from the $3$-sphere to $M,$ then the fundamental group of $M$ is finite, and by Geometrization (see [Perelman:Geometrization]) $M$ is spherical. This can also be obtained using a related result in [HughesKjuchukovaMiller:OpenBranchedCovers, Lemma 2.4]. Then the goal of the proof of Theorem 2.8 is to show the existence of universal links in spherical manifolds and the necessity condition for the non-orientable case.
Furthermore, we notice that Thurston provided a more general definition of a universal link in [Thurston:UniversalLinks], by calling a link $L$ in an arbitrary orientable closed $3$-manifold $M$ complement universal if there exists a link in any orientable closed $3$-manifold whose complement is virtually homeomorphic to $M{\smallsetminus}L$ (see Section 3). While every universal link is complement universal, the converse is not generally true, as we show:
@@
There exists a complement universal link in infinitely many $3$-manifolds which is not a universal link.
There is another homotopy obstruction, for the existence of branched covers between closed oriented manifolds, known by Hopf in [Hopf:ImageOfFiniteIndex, Satz I], which is that the induced morphism in rational cohomology algebras is injective. In this setting, Berestein and Edmons proved in [BeresteinEdmonds:BranchedCovers], among other things, that an orientable $3$-manifold $M$ is a $3$-fold branched covering of $\mathbb{S}^{2}\times\mathbb{S}^{1}$ if and only if $H^{1}(M)\neq 0$. Some time later, Núñez in [Nuñez:UniversalLinksInS2XS1, Theorem 2.1] gave an example of a link $L$ in $\mathbb{S}^{2}\times\mathbb{S}^{1}$ such that every closed, orientable $3$-manifold NN with $H^{1}(N)\neq 0$ is a branched covering over $\mathbb{S}^{2}\times\mathbb{S}^{1}$ with branching set exactly the link $L,$ and denoted these links as universal link for $\mathbb{S}^{2}\times\mathbb{S}^{1}.$
In this paper we investigate a similar problem of finding a closed, aspherical $3$-manifold $M$, and a link $L$ inside such that every closed, aspherical $3$-manifold NN is a branched covering over $M$ with branching set exactly the link $L.$ But we prove that such branched coverings can not even exist:
Proposition 4.2.
There is no closed, aspherical $3$-manifold $M$ such that every closed, aspherical $3$-manifold NN is a branched covering over $M.$
Motivated by the previous result, we investigate also the problem of finding a closed, reducible $3$-manifold $M$ such that every closed, reducible $3$-manifold NN is a branched covering over $M,$ and also proved that such branched coverings can not even exist:
@@
There is no closed reducible $3$-manifold $M$ such that every closed, reducible $3$-manifold NN is a branched covering over $M.$
More information about the existence of branched coverings between manifolds can be found in [HughesKjuchukovaMiller:OpenBranchedCovers].
The paper is organized as follows. In section 2, we give some basic definitions about branched coverings, universality and transfer maps, followed by the proof of Theorem 2.8. In section 3, we prove the non-equivalence between the definitions of complement universal links and universal links. In section 4, we address the problem of finding $3$-manifolds that could have universal links within a restricted class of closed $3$-manifolds, in Proposition 4.2, for aspherical manifolds and in Corollary 4.4, for reducible manifolds. Finally, in section 5 we present some open problems and questions concerning complement universality and characterization on the existence of branched covers between some closed $3$-manifolds.
@@ 1.1. Acknowledgements
We gratefully thank Dieter Kotschick for his useful advice and comments on an earlier preprint, mainly proving Proposition 4.2 and recalling Hopf’s paper [Hopf:ImageOfFiniteIndex]. Also, we thank Mario Eudave and Marc Kegel for pointing us toward helpful references. Finally, the authors thank CIMAT for creating an attractive mathematical environment.
@@ 2. Universal links in spherical $3$-manifolds
The following observation, originally due to Hopf, will be one of the main tools used in this work.
Proposition 2.2 (Hopf [Hopf:ImageOfFiniteIndex, Satz VIII] ).
Let $f\colon M\to N$ be a branched covering between orientable, closed, connected $n$-manifolds. Then the image of the induced homomorphism $f_{\sharp}\colon\pi_{1}(M)\to\pi_{1}(N)$ is a subgroup of finite index; that is, $\pi_{1}(N):f_{\sharp}(\pi_{1}(M))]<\infty.$
Proof.
Recall that a branched covering between closed oriented manifolds has non-zero degree; equivalently, the induced map $f_{*}\colon H_{n}(M)\to H_{n}(N)$ is non-zero.
[…]
If $\widetilde{N}$ were non-compact, then $H_{n}(\widetilde{N})=0$, implying that $f_{*}=0$. This contradicts the fact that $f$ has non-zero degree. Therefore, $\widetilde{N}$ must be compact.
Since $\widetilde{N}$ is compact, the covering map $\rho\colon\widetilde{N}\to N$ is finite, which is equivalent to $H$ being a finite-index subgroup of $\pi_{1}(N)$. This completes the proof. ∎
@@
Transfer maps. Now we explore an homological obstruction for the existence of branched covering maps between the $3$-sphere and a given $3$-manifold, which we will use to argue the non-orientable case in Theorem 2.8.
The subject of transfer maps on covering maps is well-known. But its existence can be extended to non-zero degree maps or branched covers. In [Smith:TransferRamifiedCoverings], L. Smith defined a $d$-fold ramified covering map, where branched coverings are a particular case, and proved a generalization of transfer maps for branched covers that includes non-orientable manifolds, as follows:
@@
Corollary 2.6.
Rational homology spheres do not branch cover a closed $3$-manifold $M$ with $H_{1}(M)\neq 0.$
Corollary 2.7.
Rational homology spheres do not branch cover a non-orientable, closed $3$-manifolds.
Proof.
Non-orientable closed $3$-manifolds $M$ have non-zero first Betti number [BeresteinEdmonds:BranchedCovers, Lemma 8.6]. Then by Corollary 2.6 a rational homology spheres can not branch cover a non-orientable, closed $3$-manifold. ∎
@@
A closed $3$-manifold admits a universal knot or link if and only if it is spherical.
@@
By Theorem [HughesKjuchukovaMiller:OpenBranchedCovers, Lemma 2.4] and Proposition 2.2 we have the neccessity condition when $M$ is an orientable closed $3$-manifold. For the non-orientable case, we have by Corollay 2.7 that there is not branching cover from the $3$-sphere to $M.$
Now we prove the existence of universal links. Notice that the same proof can be applied to knots.
[…]
@@ 3. Complement universal links that are not universal
Notice that if $L$ is a universal link, it is a complement universal link, because the restriction of a branched covering map on the complement of the branching set is a finite covering map (see Def. 2.3).
@@
There exists a complement universal link in infinitely many $3$-manifolds that is not universal.
@@
Consider $M$, a non-spherical $3$-manifold that is homeomorphic to a Dehn surgery over a hyperbolic universal link $L$ in a spherical $3$-manifold $S,$ e.g. three ortogonal geodesics in the $3$-torus, whose complement is homeomorphic to the Borromean ring complement in the $3$-sphere. By Thurston’s hyperbolic Dehn filling theorem, this provides infinitely many $3$-manifolds $M.$ Moreover, for every $N$ orientable closed $3$-manifold there is a branched cover $p:(N,p^{-1}(L))\rightarrow(S,L),$ which induces a finite cover from $N{\smallsetminus}p^{-1}(L)$ to $S{\smallsetminus}L,$ and the latter is homeomorphic to the complement of the core $L^{\prime}$ of the Dehn filled tori in $M.$
@@
In conclusion, $L^{\prime}\subset M$ is a complement universal link, that is not a universal link because $M$ is non-spherical by Theorem 2.8. And by varying over the integral Dehn surgery on $L$ in $S$, we obtain infinitely many non-homeomorphic $3$-manifolds $M.$
In the following section, we restrict the notion of universality to a special family of closed $3$-manifolds.
4. Non-existence of Branched covers between some $3$-manifolds
Núñez gave in [Nuñez:UniversalLinksInS2XS1] different notions of universality for links, one of which is the following:
Definition 4.1.
A universal link for $\mathbb{S}^{2}\times\mathbb{S}^{1}$is is a link $L$ in $\mathbb{S}^{2}\times\mathbb{S}^{1},$ such that every closed, orientable $3$-manifold NN with $H^{1}(N)\neq 0$ is a branched covering over $\mathbb{S}^{2}\times\mathbb{S}^{1}$ with branching set exactly the link $L.$
And gave in [Nuñez:UniversalLinksInS2XS1, Theorem 2.1] an example of a universal link for $\mathbb{S}^{2}\times\mathbb{S}^{1}.$ In the first part of this section, we prove that the analog notion of universality can not hold for the family of closed aspherical $3$-manifolds:
Proposition 4.2.
There is no closed aspherical $3$-manifold $M$ such that every closed, aspherical $3$-manifold NN is a branched covering over $M.$
Proof.
First notice that if $M$ is non-orientable, then we have a hyperbolic rational homology sphere, e.g. the Fomenko–Matveev–Weeks manifold, which can not branch cover $M,$ because of Corollary 2.7. If there is an orientable, closed, aspherical $3$-manifold $M$ such that every closed, aspherical $3$-manifold NN is a branched covering over $M.$ Then there is a non-zero degree map form $f:\mathbb{T}^{3}\rightarrow M.$ By Proposition 2.2, we have that:
  1. (1)
    $M$ is spherical, if $f_{\#}(\pi_{1}(\mathbb{T}^{3}))=\{e\},$ or
  2. (2)
    $M$ is virtually abelian, meaning $M$ has a finite covering with fundamental group $\mathbb{Z}^{k},$ if $f_{\#}(\pi_{1}(\mathbb{T}^{3}))\neq\{e\}.$
The first case is impossible, because $M$ is aspherical. And in the second case, again by asphericity we have that $M$ is finitely covered by the 3-torus. However, there are closed aspherical $3$-manifolds, e.g. the ones with $Nil$-geometry, that are not branched covers over $M$ because of [KotschickNeofytidis:DominationByCircleBunles, Theorem 2 and Theorem 6]. ∎
In the second part of this section, we prove in Theorem 4.3 that the only closed, reducible $3$-manifolds, to be branch covered by $P^{3}\#P^{3},$ is itself. And conclude in Corollary 4.4 that the analog notion of universality in Definition 4.1 can not hold for the family of closed, reducible $3$-manifolds.
Theorem 4.3.
Let $M$ be a closed $3$-manifold. There exists a branched covering
$f\colon P^{3}\#P^{3}\longrightarrow M$
if and only if
  1. (1)
    $M$ is a spherical manifold, or
  2. (2)
    $M$ is homeomorphic to $P^{3}\#P^{3}$.
Proof.
The converse is clearly true, because as we have shown in Theorem 2.8 all spherical $3$-manifolds are branch covered by any orientable, closed $3$-manifold. And when $M=P^{3}\#P^{3}$ we have the identity map $P^{3}\#P^{3}\to P^{3}\#P^{3}$ as a branched covering.
Suppose there exists a branched covering $f\colon P^{3}\#P^{3}\to M$. By Corollary 2.7, the manifold $M$ must be orientable. From now on, we assume that $M$ is orientable; in particular, $f$ is a map of non-zero degree.
Let $M=M_{1}\#\cdots\#M_{k}$ be the (unique) prime decomposition of $M$, given by the Kneser–Milnor prime decomposition theorem. Each prime summand $M_{i}$ is either aspherical, $\mathbb{S}^{1}\times\mathbb{S}^{2}$, or spherical.
Observe that no summand $M_{i}$ can be $\mathbb{S}^{1}\times\mathbb{S}^{2}$. Indeed, if this were the case, collapsing all other summands and composing with $f$ would yield a non-zero degree map $P^{3}\#P^{3}\to\mathbb{S}^{1}\times\mathbb{S}^{2},$ which is impossible by Corollary 2.6. Hence, each summand $M_{i}$ is either aspherical or spherical.
For each $i$, consider the map $f_{i}\colon P^{3}\#P^{3}\to M_{i}$ obtained by composing $f$ with the non-zero degree map $M\to M_{i}$ that collapses all connected summands other than $M_{i}$.
Suppose that $M_{i}$ is aspherical. Then $\pi_{1}(M_{i})$ has no elements of order two (it does not have torsion elements in general). Since $\pi_{1}(P^{3}\#P^{3})\cong\mathbb{Z}_{2}*\mathbb{Z}_{2}$, it follows that the induced homomorphism $(f_{i})_{\#}\colon\pi_{1}(P^{3}\#P^{3})\to\pi_{1}(M_{i})$ is trivial, that is, $(f_{i})_{\#}(\pi_{1}(P^{3}\#P^{3}))=\{e\}.$
By Proposition 2.2, this implies that $M_{i}$ is spherical, contradicting the assumption that $M_{i}$ was aspherical. Therefore, all summands $M_{i}$ must be spherical.
Again by Proposition 2.2, the image $f_{\#}(\pi_{1}(P^{3}\#P^{3}))$ is a subgroup of finite index in $\pi_{1}(M)=*_{i}\pi_{1}(M_{i}).$ By the Kurosh subgroup theorem [kurosch], the images of the generators of
$\pi_{1}(P^{3}\#P^{3})\cong\langle x,y\mid x^{2}=y^{2}=e\rangle$
must be conjugate into the free factors $\pi_{1}(M_{i})$. That is, there exist elements $w_{1},w_{2}\in\pi_{1}(M)$ and indices $i,j$ such that
$w_{1}f_{\#}(x)w_{1}^{-1}\in\pi_{1}(M_{i})\quad\text{and}\quad w_{2}f_{\#}(y)w_{2}^{-1}\in\pi_{1}(M_{j}).$
There are three possibilities: both, one, or neither of the elements $f_{\#}(x)$ and $f_{\#}(y)$ is trivial. If at least one of them is trivial, then the image $f_{\#}(\pi_{1}(P^{3}\#P^{3}))$ is finite. By Proposition 2.2, this implies that $M$ is spherical.
Thus, we are left with the case in which neither $f_{\#}(x)$ nor $f_{\#}(y)$ is trivial. In this situation, again by the Kurosh subgroup theorem, the image is isomorphic to $H=\mathbb{Z}_{2}*\mathbb{Z}_{2}.$
By Proposition 2.2, $H$ must be a finite-index subgroup of $G=*_{i}A_{i}$, where $A_{i}=\pi_{1}(M_{i})$ are finite groups.
Using the Euler characteristic of groups (see [Toshiyuki:EulerCaracteristicOfGroups]), we have
$\chi(H)=[G:H]\,\chi(G).$
Since $\chi(H)=\chi(\mathbb{Z}_{2})+\chi(\mathbb{Z}_{2})-1=0,$ it follows that $\chi(G)=0$. On the other hand,
$\chi(G)=\left(\sum_{i=1}^{k}\frac{1}{|A_{i}|}\right)-k+1.$
This equality can hold only when $k=2$ and $|A_{1}|=|A_{2}|=2$.
Therefore, in this final case, the manifold $M$ has exactly two spherical summands with fundamental group $\mathbb{Z}_{2}$, and hence $M\cong P^{3}\#P^{3}$. This completes the proof.
Corollary 4.4.
There is no closed reducible $3$-manifold $M$ such that every closed, reducible $3$-manifold NN is a branched covering over $M.$
Proof.
If there is an orientable, closed, reducible $3$-manifold $M$ such that every closed, reducible $3$-manifold NN is a branched covering over $M.$ Then by Theorem 4.3, $M=P^{3}\#P^{3}.$
Furthermore, if we choose $p\in\mathbb{Z}_{>2},$ then we would have a non-zero degree map from the connected sum of two lens spaces of order $p$ to $P^{3}\#P^{3}.$ But we have a contradiction, because due to Proposition 2.2, $P^{3}\#P^{3}$ would be spherical. ∎